# Flat band mediated photon-photon interactions in 2D waveguide QED networks

### v2

@@ I Introduction
Systems featuring dispersionless (flat) bands provide an ideal setting for the emergence of strong particle correlations, as interactions dominate over single-particle energy scales [1]. Besides the famous Landau level problem, which gives rise to fractional quantum Hall states [2], and the recent discovery of strongly correlated states in twisted bilayer graphene [3, 4], flat bands can also arise in a variety of lattice models [5, 6]. By definition, single particles in a flat band exhibit zero group velocity and infinite effective mass. While in trivial flat bands, this comes from having decoupled sites, in flat bands embedded within multi-band systems, it results from destructive wave interference, leading to the formation of highly localized states. These states, known as compact localized states (CLS), are superpositions of Bloch states confined to a few lattice sites [7, 8, 9, 10], which can even result in the extreme situation of full localization with all-bands flat known as Aharonov-Bohm caging [11, 12]. In the multi-particles case, interactions can induce transport that occurs in pairs [13, 14, 15, 16, 17], something reminiscent of Cooper pairing in superconductors, that leads to superfluid behavior [18, 19, 20]. Remarkably, the effective mass of these pairs has been shown to have a geometric origin based on the quantum metric of the flat band [21, 22]. These fascinating effects of flat-band physics have been extensively studied and experimentally probed [23, 24, 25, 26, 27, 28, 29] in models with finite-range hopping. In contrast, long-range hopping has been identified as a mean to introduce non-trivial topology to flat bands (non-zero Chern number)[30, 31]. However, it is generally understood that incorporating fine-tuned long-range terms for topological bands breaks the single-particle localization due to the obstruction theorem and the non-zero Chern number, thus inducing a finite, possibly small, band dispersion, making the described effects less robust [32].
@@
Figure 1: (a) A Lieb lattice of quantum emitters with lattice constant $d$ and intracell distance $a$ is coupled to a square grid of one-dimensional waveguides. The unit cell of the lattice consists of a corner site and two edge sites coupled to two waveguides and an individual waveguide, respectively. The quantum emitters are modeled as nonlinear resonators with on-site repulsive interaction $U$. (b) The single-excitation dispersion within the first Brillouin zone: two dispersion branches are separated by an isolated flat band at zero energy. Here, we set $k_{0}d=\pi$ and $a=\frac{d}{2}$.
@@
Infinite long-range photon-mediated coupling arise naturally among multiple emitters coupled to a one-dimensional photonic channel. This scenario, referred to as waveguide QED [33, 34, 35], has attracted increasing attention due to its feasibility for experimental implementation either at optical [36, 37, 38, 39, 40] or microwave frequencies [41, 42, 43, 44, 45, 46, 47]. Extensive studies have been conducted to tailor atom-atom interactions through photonic bath engineering. In particular, when the atomic frequency falls within a photonic bandgap, it becomes possible to induce coherent atom-atom interactions with tunable ranges [48, 49, 50, 51]. Further control over these interactions can be achieved by coupling the atoms to topological photonic lattices [52, 53, 54, 55] or photonic flat bands [56, 57, 58]. A fundamentally different approach to engineering coherent and dissipative interaction properties involves leveraging the collective emission of ordered atomic arrays coupled to propagating photons. In this scenario, collective phenomena can significantly alter the excitation lifetime within the array, leading to either superradiant [59, 60, 61] or subradiant states [62, 63, 64, 65, 66, 67, 68, 69, 70]. Such collective interactions also influence the coherent properties of the emitters, giving rise to polariton-like dispersion branches [63, 71, 72], which determine the propagation speed of excitations within the array. The engineering of these dispersions has been studied at the single-excitation level in two dimensions [73, 74, 75, 76, 77, 78], uncovering the presence of topological [79, 80, 81] and flat bands [81, 82, 83]. However, in the latter case, no demonstration of compact localized state formation or investigation of multi-excitation physics was conducted. Meanwhile, the coupling of quantum emitters with confined light is also known to induce effective photon-photon interactions, leading to highly correlated light observable as photon anti-bunching [84, 85, 86] and bunching [87, 88, 40, 89, 90]. These phenomena, involving single or multiple emitters, have been extensively explored in one-dimensional channels  [91, 92, 93, 94, 71, 95], as well as in arrays of emitters coupled to waveguide networks [96, 97], or to two-dimensional waveguides [98].
@@ II Model
We consider a two-dimensional quantum network consisting of an array of $N$ quantum emitters (atoms) coupled to a square grid of one-dimensional, disconnected photonic waveguides, as illustrated in Fig. 1. We describe the quantum emitters as bosons represented by the field operator $\hat{b}_{i}$ ($\hat{b}_{i}^{\dagger}$) annihilating (creating) a photon at position $i$ and satisfying $[\hat{b}_{i},\hat{b}_{j}^{\dagger}]=\delta_{ij}$. To make energetically unfavorable for an emitter to host multiple excitations, we include an onsite photon-photon repulsion of strength $U$, that describes a photonic nonlinearity [99]. In the limit of infinite interaction strength $U$, the emitters effectively behave as two-level systems (qubits) with ground and excited states $|g\rangle,|e\rangle$ with corresponding transition frequency $\omega_{eg}$. The emitters are arranged in a Lieb lattice [5], composed of: two edge sites, A and C, each coupled via the $|g\rangle-|e\rangle$ transition exclusively to one vertical or horizontal waveguide, and one corner site, B, coupled via the same transition to two waveguides at their crossing point. We denote the lattice constant by $d$, while the intracell distance, assumed to be the same in both directions (A-B and B-C), is indicated by $a$.
The one-dimensional waveguides mediate long-range interactions among the emitters according the network connectivity. Assuming a linear dispersion relation for the photons, $\omega_{k}=c|k|$, where $c$ is the photon group velocity and $k$ is the wavevector, and neglecting time-delay effects, we apply a Born-Markov approximation to adiabatically eliminate the photonic degrees of freedom. The resulting effective photon-mediated dissipative and coherent interactions are encoded in the non-Hermitian Hamiltonian (with $\hbar=1$)  [33, 100, 101, 102, 103]
@@
where $\mathbf{x}_{i}$ represents the positions of the atoms in the lattice, $\gamma$ is the decay rate of an individual emitter coupled to a 1D waveguide, and $k_{0}=\omega_{eg}/c$ is the photon wavevector resonant with the $|g\rangle\leftrightarrow|e\rangle$ transition. The sum $(i,j)\in\mathcal{N}$ runs only over emitter pairs connected by the waveguide network $\mathcal{N}$, as shown in Fig. 1. Notice that, as each waveguide is treated as an independent photonic bath, the diagonal terms $i=j$, associated with individual spontaneous emission processes, are characterized by a rate $\gamma$ for the edge emitters A and C, and a rate $2\gamma$ for the corner sites B.
Hamiltonian (1) fully describes the system within a fixed excitation sector, provided that external pumping fields are not included in the model. In the following section, we focus on the single-excitation sector and later on we analyze the two-excitation sector.
III Flat band and compact localised states
@@ III.1 Single particle dispersion
where $\epsilon^{(1)}_{1D}(k_{i})$ is the single excitation dispersion of a 1D atomic array coupled to a one-dimensional waveguide [63],
$\epsilon_{1D}^{(1)}(k_{i})$ $=\frac{\gamma}{4}\left[\mathrm{ctg}\left(\frac{(k_{0}+k_{i})d}{2}\right)+\mathrm{ctg}\left(\frac{(k_{0}-k_{i})d}{2}\right)\right]\,,$ (4)
while the off-diagonal terms $t(k_{i})$ read:
$\begin{split}&t(k_{i})=\epsilon_{1D}^{(1)}(k_{i})\cos(k_{0}a)+\frac{\gamma}{2}\sin(k_{0}a)\\ &-i\frac{\gamma}{4}\sin(k_{0}a)\left[\mathrm{ctg}\left(\frac{(k_{0}+k_{i})d}{2}\right)-\mathrm{ctg}\left(\frac{(k_{0}-k_{i})d}{2}\right)\right]\,.\end{split}$ (5)
@@
$\begin{split}&\lambda_{1}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\\ \end{pmatrix},\quad\lambda_{2}=\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\\ \end{pmatrix},\\ &\lambda_{4}=\begin{pmatrix}0&0&1\\ 0&0&0\\ 1&0&0\\ \end{pmatrix},\quad\lambda_{5}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\\ \end{pmatrix},\end{split}$ (7)
@@
^$\begin{split}\hat{\mathcal{H}}(\mathbf{k})&=\frac{\gamma}{2}\sin(k_{0}a)\Big{[}(\lambda_{1}+\lambda_{4})\\ &-\tan\left(\frac{k_{x}d}{2}\right)\lambda_{2}-\tan\left(\frac{k_{y}d}{2}\right)\lambda_{5}\Big{]}\,.\end{split}$ (8)
By diagonalizing the Bloch Hamiltonian in Eq. (8), we obtain the single-excitation spectrum, $\epsilon^{(1)}(\mathbf{k})$, shown in Fig. 1(b). This spectrum consists of two polariton-like dispersive bands separated by a flat band with $\epsilon^{(1)}(\mathbf{k})=0$. The presence of the gap at the $\Gamma$ point is due to the $(\lambda_{1}+\lambda_{4})$-term in the Hamiltonian ^$\hat{\mathcal{H}}(\mathbf{k})$. In the standard nearest-neighbor Lieb lattice model, this term appears, for example, when breaking inversion symmetry, while here originates from the long-range structure of the couplings. The two dispersive bands are separated from the flat-band by an energy gap $\Delta\epsilon=\frac{\gamma}{\sqrt{2}}|\sin(k_{0}a)|$. This gap is crucial for defining a complete basis of compact localized states (CLS), as we discuss in the following. A large energy gap is also essential for observing flat-band-mediated photon-photon interactions, which will be explored in the next section. Therefore, throughout the paper, we maximize this gap to $\Delta\epsilon=\gamma/\sqrt{2}$ by setting $a=d/2$. Finally, note that the dispersive bands diverge at the edge of the Brillouin zone, as typically occurs in 1D wQED systems [33]. These divergences are associated with superradiant modes at the resonant wavevector $|\mathbf{k}|=k_{0}$ [63, 104, 105].

### v3

@@ I Introduction
Systems featuring dispersionless (flat) bands provide an ideal setting for the emergence of strong particle correlations, as interactions dominate over single-particle energy scales [1]. Besides the famous Landau level problem, which gives rise to fractional quantum Hall states [2], and the recent discovery of strongly correlated states in twisted bilayer graphene [3, 4], flat bands can also arise in a variety of lattice models [5, 6]. By definition, single particles in a flat band exhibit zero group velocity and infinite effective mass. While in trivial flat bands, this comes from having decoupled sites, in flat bands embedded within multi-band systems, it results from destructive wave interference, leading to the formation of highly localized states. These states, known as compact localized states (CLS), are superpositions of Bloch states confined to a few lattice sites [7, 8, 9, 10], which can even result in the extreme situation of full localization with all-bands flat known as Aharonov-Bohm caging [11, 12]. In the multi-particles case, interactions can induce transport that occurs in pairs [13, 14, 15, 16, 17, 18], something reminiscent of Cooper pairing in superconductors, that leads to superfluid behavior [19, 20, 21]. Remarkably, the effective mass of these pairs has been shown to have a geometric origin based on the quantum metric of the flat band [22, 23]. These fascinating effects of flat-band physics have been extensively studied and experimentally probed [24, 25, 26, 27, 28, 29, 30] in models with finite-range hopping. In contrast, long-range hopping has been identified as a mean to introduce non-trivial topology to flat bands (non-zero Chern number)[31, 32]. However, it is generally understood that incorporating fine-tuned long-range terms for topological bands breaks the single-particle localization due to the obstruction theorem and the non-zero Chern number, thus inducing a finite, possibly small, band dispersion, making the described effects less robust [33].
@@
Figure 1: (a) A Lieb lattice of $N$ quantum emitters with lattice constant $d$ and intracell distance $a$ is coupled to a square grid of one-dimensional waveguides. The unit cell of the lattice consists of a corner site and two edge sites coupled to two waveguides and an individual waveguide, respectively. The quantum emitters are modeled as nonlinear resonators with on-site repulsive interaction $U$. (b) The single-excitation dispersion within the first Brillouin zone: two dispersion branches are separated by an isolated flat band at zero energy. Here, we set $k_{0}d=\pi$ and $a=\frac{d}{2}$.
@@
Infinitely long-range photon-mediated interactions arise naturally among multiple emitters coupled to a one-dimensional photonic channel. This scenario, referred to as waveguide QED [34, 35, 36], has attracted increasing attention due to its feasibility for experimental implementation either at optical [37, 38, 39, 40, 41] or microwave frequencies [42, 43, 44, 45, 46, 47, 48]. Extensive studies have been conducted to tailor atom-atom interactions through photonic bath engineering. In particular, when the atomic frequency falls within a photonic bandgap, it becomes possible to induce coherent atom-atom interactions with tunable ranges [49, 50, 51, 52]. Further control over these interactions can be achieved by coupling the atoms to topological photonic lattices [53, 54, 55, 56] or photonic flat bands [57, 58, 59]. A fundamentally different approach to engineering coherent and dissipative interaction properties involves leveraging the collective emission of ordered atomic arrays coupled to propagating photons. In this scenario, collective phenomena can significantly alter the excitation lifetime within the array, leading to either superradiant [60, 61, 62] or subradiant states [63, 64, dimer_shermet, 40, 65, 66, 67, 68, 69]. Such collective interactions also influence the coherent properties of the emitters, giving rise to polariton-like dispersion branches [64, 70, 71], which determine the propagation speed of excitations within the array. The engineering of these dispersions has been studied at the single-excitation level in two dimensions [72, 73, 74, 75, 76, 77], uncovering the presence of topological [78, 79, 80] and flat bands [80, 81, 82]. However, in the latter case, no demonstration of compact localized state formation or investigation of multi-excitation physics was conducted. Meanwhile, the coupling of quantum emitters with confined light is also known to induce effective photon-photon interactions, leading to highly correlated light observable as photon anti-bunching [83, 84, 71] and bunching [85, 86, 41, 87, 88]. These phenomena, involving single or multiple emitters, have been extensively explored in one-dimensional channels [89, 90, 91, 92, 70, 93], as well as in arrays of emitters coupled to waveguide networks [94, 95], or to two-dimensional waveguides [96].
@@ II Model
We consider an array of $N$ quantum emitters (atoms) coupled to a two-dimensional waveguide network. We describe the quantum emitters as bosons represented by the field operator $\hat{b}_{i}$ ($\hat{b}_{i}^{\dagger}$) annihilating (creating) a photon at position $i$ and satisfying $\hat{b}_{i},\hat{b}_{j}^{\dagger}]=\delta_{ij}$. To make energetically unfavorable for an emitter to host multiple excitations, we include an onsite photon-photon repulsion of strength $U$, that describes a photonic nonlinearity [97]. In the limit of infinite interaction strength $U$, the emitters effectively behave as two-level systems (qubits) with ground and excited states $|g\rangle,|e\rangle$ with corresponding transition frequency $\omega_{eg}$.
Before presenting the waveguide network considered in this work, we first review the derivation of the effective waveguide QED Hamiltonian for an array of quantum emitters coupled to a one dimensional photonic channel, starting from the original atom–light interaction model.
II.1 Waveguide QED Hamiltonian
Let us assume the emitters to be arranged in an ordered array with lattice constant $a$, each coupled with the same strength to a one-dimensional waveguide through the $|g\rangle\leftrightarrow|e\rangle$ transition, resulting in an individual decay rate $\gamma$. Within the frequency range relevant to the system dynamics, set by $\gamma$, we assume a linear dispersion relation for the electromagnetic field in the waveguide, $\omega_{k}=c|k|$. We can thus separately treat left and right propagating photons traveling with opposite speed $\pm c$. The Hamiltonian that describes this system under the rotating-wave and dipole approximations reads ($\hbar=1$) [34]:
$\begin{split}&\hat{H}=\hat{H}_{0}+\hat{H}_{\rm f}+\hat{H}_{\rm int}\,,\\ &\hat{H}_{0}=\sum_{i}\omega_{eg}\hat{b}_{i}^{\dagger}\hat{b}_{i}+\frac{U}{2}\sum_{i}\hat{b}_{i}^{\dagger}\hat{b}_{i}^{\dagger}\hat{b}_{i}\hat{b}_{i}\,,\\ &\hat{H}_{\rm f}=\sum_{\lambda=R,L}\int d\omega\,\omega\,\hat{a}^{\dagger}_{\lambda}(\omega)\hat{a}_{\lambda}(\omega)\,,\\ &\hat{H}_{\rm int}=\sum_{\lambda=R,L}\sum_{i}\sqrt{\frac{\gamma}{2\pi}}\int d\omega\,\omega\,\hat{a}^{\dagger}_{\lambda}(\omega)\,\hat{b}_{i}\,e^{\mp i\omega x_{i}/c}+{\rm H.c.}\end{split}$ (1)
where the bosonic operators $\hat{a}_{\lambda}(\omega)$ annihilate right- ($\lambda=R$) and left- ($\lambda=L$) propagating photons of frequency $\omega$ and wavevector $k=\pm\omega/c$, while $x_{i}$ denotes the position of the $i$-th emitter along the waveguide. This Hamiltonian describes the exchange of photons between the emitters mediated by the waveguide.
To make the analysis of this complex interacting system more tractable, we assume that the usual requirements for the Born–Markov approximation are fulfilled. This holds for the linear dispersion relation considered, which yields a smooth density of states, provided that retardation effects due to the finite speed of light $c$ can be neglected. The condition may be expressed by demanding that the characteristic timescale of atom–photon dynamics, set by the spontaneous emission rate $\gamma$ of a single emitter into the bath, is much longer than the photon propagation time across the entire system, i.e., $\gamma\ll c/(Nd)$. Within this regime, photon-mediated interactions between emitters can be treated as instantaneous and the photonic modes can be adiabatically eliminated, leading to the following Lindblad master equation governing the dynamics of the emitters density operator $\hat{\rho}(t)$ [34, 98, 99, 100]:
[…]
where
$\hat{H}_{S}=\frac{\gamma}{2}\sum_{i,j}\sin{\left(k_{0}|x_{i}-x_{j}|\right)}\,\hat{b}_{i}^{\dagger}\hat{b}_{j},$ (3)
describes the coherent collective dynamics of the emitters and $k_{0}=\omega_{eg}/c$ is the photon wavevector resonant with the $|g\rangle\leftrightarrow|e\rangle$ transition. The second term in Eq. (2) describes collective dissipation governed by the Hermitian dissipation matrix:
$\Gamma_{ij}=\frac{\gamma}{2}\cos(k_{0}|x_{i}-x_{j}|).$ (4)
To proceed, we rewrite the Lindblad master equation in Eq. (2) in the form
$\frac{d\hat{\rho}}{dt}=-i\Big[\hat{H}_{\rm eff}\hat{\rho}-\hat{\rho}\hat{H}_{\rm eff}^{\dagger}\Big]+2\sum_{i,j}\Gamma_{ij}\hat{b}_{i}\hat{\rho}\hat{b}_{j}^{\dagger}\,$, (5)
where the anti-commutator part of the master equation has been included into the waveguide QED non-Hermitian Hamiltonian, defined as [34]:
$\hat{H}_{\rm eff}=-i\frac{\gamma}{2}\sum_{i,j}e^{-ik_{0}|x_{i}-x_{j}|}\hat{b}_{i}^{\dagger}\hat{b}_{j}+\frac{U}{2}\sum_{i}\hat{b}_{i}^{\dagger}\hat{b}_{i}^{\dagger}\hat{b}_{i}\hat{b}_{i}\,$, (6)
where the last term accounts for the emitters nonlinearity.
This effective Hamiltonian fully describes photon-mediated interactions among the emitters within a fixed excitation sector, provided that external pumping fields are not included in the model. For each excitation sector the eigenvalues are of the form
$E_{1D}^{(n_{e})}(k)=\epsilon_{1D}^{(n_{e})}(k)-i\gamma_{1D}^{(n_{e})}(k)/2,$ (7)
where $n_{e}$ is the number of excitations and $\epsilon^{(n_{e})}_{s}$ and $\gamma^{(n_{e})}_{s}$ denote the energy and collective decay rate of the $k$-th state, respectively. In the single-excitation sector, the energy spectrum defines the 1D wQED dispersion relation [64]
$\epsilon_{1D}^{(1)}(k_{i})$ $=\frac{\gamma}{4}\left[\mathrm{\cot}\left(\frac{(k_{0}+k_{i})d}{2}\right)+\mathrm{\cot}\left(\frac{(k_{0}-k_{i})d}{2}\right)\right]\,$, (8)
while the decay rates are suppressed as $N^{-3}$ with increasing system size and vanish in the thermodynamic limit, where the emitter array has no edges [64]. Consequently, the effective Hamiltonian (6), projected onto this excitation sector, becomes Hermitian in this limit. For higher excitations, this remains true, albeit with a different scaling of the decay rates [91, 65], up to the limiting case of half-filling [69].
II.2 Two-dimensional waveguide QED network
We now extend the previously developed formalism to treat an array of quantum emitters coupled to a square grid of one-dimensional, disconnected photonic waveguides, as illustrated in Fig. 1. The emitters are arranged in a Lieb lattice [5], composed of two edge sites, A and C, each coupled via the $|g\rangle-|e\rangle$ transition exclusively to one vertical or horizontal waveguide, and one corner site, B, coupled via the same transition to two waveguides at their crossing point. We denote the lattice constant by $d$, while the intracell distance, assumed to be the same in both directions (A-B and B-C), is indicated by $a$. The one-dimensional waveguides mediate long-range interactions among the emitters according to the network connectivity. Since the waveguides are independent, the master equation derived for a single waveguide (5) can be straightforwardly generalized to the entire waveguide network, with each waveguide providing an additive contribution [101]. The resulting non-Hermitian Hamiltonian then reads:
@@ II.2 Two-dimensional waveguide QED network
where $\mathbf{x}_{i}$ represents the positions of the atoms in the lattice. The sum $(i,j)\in\mathcal{N}$ runs only over emitter pairs connected by the waveguide network $\mathcal{N}$, as shown in Fig. 1. Notice that, as each waveguide is treated as an independent photonic bath, the diagonal terms $i=j$, associated with individual spontaneous emission processes, are characterized by a rate $\gamma$ for the edge emitters A and C, and a rate $2\gamma$ for the corner sites B. In the following analysis, we will consider an array in the plane with open boundary conditions and $N_{c}=N/3$ unit cells.
III Flat band and compact localized states
@@ III.1 Single particle dispersion
where $\epsilon^{(1)}_{1D}(k_{i})$, with $i=x,y$, is the single excitation dispersion of a 1D atomic array coupled to a one-dimensional waveguide given in Eq. (8), while the off-diagonal terms $t(k_{i})$ read:
$\begin{split}&t(k_{i})=\epsilon_{1D}^{(1)}(k_{i})\cos(k_{0}a)+\frac{\gamma}{2}\sin(k_{0}a)\\ &-i\frac{\gamma}{4}\sin(k_{0}a)\left[\mathrm{\cot}\left(\frac{(k_{0}+k_{i})d}{2}\right)-\mathrm{\cot}\left(\frac{(k_{0}-k_{i})d}{2}\right)\right]\,.\end{split}$ (12)
@@
[…]
@@
$\begin{split}\hat{\mathcal{H}}(\mathbf{k})&=\frac{\gamma}{2}\sin(k_{0}a)\Big[(\lambda_{1}+\lambda_{4})\\ &-\tan\left(\frac{k_{x}d}{2}\right)\lambda_{2}-\tan\left(\frac{k_{y}d}{2}\right)\lambda_{5}\Big]\,.\end{split}$ (15)
By diagonalizing the Bloch Hamiltonian in Eq. (15), we obtain the single-excitation spectrum, $\epsilon^{(1)}(\mathbf{k})$, shown in Fig. 1(b). This spectrum consists of two polariton-like dispersive bands separated by a flat band with $\epsilon^{(1)}(\mathbf{k})=0$. The presence of the gap at the $\Gamma$ point is due to the $(\lambda_{1}+\lambda_{4})$-term in the Hamiltonian $\hat{\mathcal{H}}(\mathbf{k})$. In the standard nearest-neighbor Lieb lattice model, this term appears, for example, when breaking inversion symmetry, while here originates from the long-range structure of the couplings. The two dispersive bands are separated from the flat-band by an energy gap $\Delta\epsilon=\frac{\gamma}{\sqrt{2}}|\sin(k_{0}a)|$. This gap is crucial for defining a complete basis of compact localized states (CLS), as we discuss in the following. A large energy gap is also essential for observing flat-band-mediated photon-photon interactions, which will be explored in the next section. Therefore, throughout the paper, we maximize this gap to $\Delta\epsilon=\gamma/\sqrt{2}$ by setting $a=d/2$. Finally, note that the dispersive bands diverge at the edge of the Brillouin zone, as typically occurs in 1D wQED systems [34]. These divergences are associated with superradiant modes at the resonant wavevector $|\mathbf{k}|=k_{0}$ [64, 102, 103].